How many time constants must elapse for initially uncharged capacitor in an RC series circuit to be charged to 99% of its equilibrium charge?

Grade:12

1 Answers

Saurabh Kumar
9 years ago
Charge q across the capacitor is q = q0(1 − e-t/T)
Solve 0.99 = (1- e-t/T) to find t/T = 0.46

or t = 0.46 T